7. Among 13,963,753 males and 14,272,325 females, there were 20,734 cases of prostate cancer and 19,107 cases of female breast cancer respectively. Calculate the incidence rates of prostate cancer and female breast cancer.

Express your answers using an appropriate rate base. [5]
Answer #1

incidence rate of prostate cancer be :-

number of cases of prostate cancer total number of males [ as prostate cancer happens in male only]

20734 13963753

1485 + 10-3

= 148.5*10=-

= 148.5 per 100000

incidence rate of female breast cancer be :-

number of cases of female breat cancer total number of females

19107 14272325

<1.339 + 10-3

= 133.9 + 105

= 133.9 per 100000

*** if you have any doubt regarding the problem please write it in the comment box.if you are satisfied please give me a LIKE if possible..